Review ▼ Part A A 4.5 cm × 4.4 cm rectangle lies in the xy-plane What is the electric flux through the rectangle if E = (100 i-200k) N/C? You may want to review (Pages 664 -668) Submit Request Answer
Answer #1

Given that Answer: - A = (4.5 cm x Gele Com) E = (19.8 104 m2) E E = (1008 - 2008) / part (A) : P ER= (160 2 - 2004) . (19.8
Answer #2

The Area vector is in the y-direction, not the z-direction, so the electric flux should be 0.

ΦE = E●A = (100i + 0j - 200k) ● (0i + 1.75*10^-3 j +0k)

      = (100*0) + (0*1.75*10^-3) + (-200*0)

      = 0

source: Pearson
answered by: Cindy